OpenStudy (anonymous):

show that the lim |xn+1-xn|=0

OpenStudy (unklerhaukus):

\[\lim_{n\to?}|x^{n+1}-x^n|=0\]

OpenStudy (anonymous):

they're sequences

OpenStudy (unklerhaukus):

what is it the limit of? n approaches ?

OpenStudy (anonymous):

it doesn't say in the question, I'll assume infinity

OpenStudy (anonymous):

can you post a screen shot? this question is totally incomplete

OpenStudy (anonymous):

ok give me a minute

OpenStudy (anonymous):

oh wait I forgot it was part of a question

ganeshie8 (ganeshie8):

\[\lim \left|x_{n+1}-x_n\right| = \lim \left|\sum\limits_{k=1}^{n+1}\frac{1}{k} - \sum\limits_{k=1}^n\frac{1}{k}\right|\]

ganeshie8 (ganeshie8):

split the first sum and cancel it with the second sum

OpenStudy (anonymous):

Okay dumb moment.. the first sum wouldn't be 1/2-1?

ganeshie8 (ganeshie8):

we don't need to calculate the sum just notice that \(\large x_{n+1} - x_n\) gives you the (n+1)th term of the series

ganeshie8 (ganeshie8):

\[\begin{align} \lim \left|x_{n+1}-x_n\right| &= \lim \left|\sum\limits_{k=1}^{n+1}\frac{1}{k} - \sum\limits_{k=1}^n\frac{1}{k}\right| \\~\\ &=\lim \left|\frac{1}{n+1}+\sum\limits_{k=1}^{n}\frac{1}{k} - \sum\limits_{k=1}^n\frac{1}{k}\right| \\~\\ &=\lim \left|\frac{1}{n+1}\right| \\~\\ \end{align}\]

ganeshie8 (ganeshie8):

take the limit

OpenStudy (anonymous):

Ahhhhh, okay perfect! thanks!

ganeshie8 (ganeshie8):

lim is interpreted as limit of the sequence it is assumed that n tends to +infinity